# Influence of momentum confinement in the phase-space dynamics of the Kramers-Henneberger atom

### v1

@@ Abstract
We investigate the phase-space dynamics of the Kramers Henneberger (KH) atom solving the time-dependent Schrödinger equation for reduced-dimensionality models and using Wigner quasiprobability distributions. We find that, for the time-averaged KH potential, coherent superpositions of eigenstates perform a cyclic motion confined in momentum space, whose frequency is proportional to the energy difference between the two KH eigenstates. This cyclic motion is also present if the full time dependent dynamics are taken into consideration. However, there are time delays regarding the time-averaged potential, and some tail-shaped spilling of the quasiprobability flow towards higher momentum regions. These tails are signatures of ionization, indicating that vanishingly small drift momentum is instrumental for stabilization. A comparison of the quasiprobability flow with classical phase-space constraints shows that, for the KH atom, confinement occurs in momentum space. This is a major difference from a molecule, for which the quasiprobability flow is confined in position space for small internuclear separation. Furthermore, we assess the stability of different propagation strategies and find that the most stable scenario for the full dynamics is obtained if the system is initially prepared in the KH ground state.
@@ I Introduction
Thus, a legitimate question is how far the analogy between the KH atom and a molecule can be pushed, and how stable the light-induced potentials are. To perform these studies, it is convenient to write the full Hamiltonian in an oscillatory frame of reference, which quivers with the electron in the presence of the laser field: the KH frame. In the KH frame, the time dependence of the field is embedded in the binding potential. Thereby, an important indicator is how far the system’s dynamics is from that obtained for the time-averaged KH potential. This time-averaged potential is the zeroth term of the Fourier expansion of the time-dependent potential and has a dichotomous structure and eigenstates of well-defined shapes and parities. A stable population in the KH eigenstates is evidence that the stabilization regime has been reached [19, 20], as well as the electronic wave packet developing a double-peaked structure. Early studies using Floquet theory also indicate that the Floquet quasienergy states exhibit a structure resembling those of the time-averaged KH potential [21]. Different coherent superpositions of KH eigenstates can be obtained, for instance, by employing fields of different lengths [22].
@@
Important tools for understanding a system’s dynamics are provided by the phase space (for a review see [23]). Besides providing information about the spatial localization of a time-dependent wave packet, phase-space tools also shed light on how its momentum distribution evolves in time, and whether its temporal evolution exhibits nonclassical features. For instance, subtleties in the dynamics of strong-field ionization in atoms [24, 25, 26] and molecules [27, 28, 29] or high-order harmonic generation in inhomogeneous fields [30], and rescattering [31, 32, 33], and how quantum effects facilitate enhanced ionization [29, 23] have been understood with the help of the phase space. Furthermore, Wigner quasiprobability distributions have been employed to characterize non-classical light [34], and in laser-induced core dynamics [35, 36] and attosecond interferometry [37].
@@
Within strong-field atomic stabilization, quantum phase-space tools have been employed in early work, following the line of inquiry of whether stabilization was classical or quantum [21, 38, 39, 22], and to understand the signatures of wave packets driven by different field types [22]. Tails in the Wigner quasi probability distributions were associated with ionization and a dichotomous structure was linked to the KH potential [39, 22]. Another area of investigation was the nonclassicality of time-dependent terms beyond the zeroth order, associating multiphoton transitions with the corrections to the KH potential [38], time averaged potentials and scarring in the phase-space distributions [21]. In addition to those, there are also studies using the classical phase space, employing periodic orbit theory and Poincaré sections [40, 41]. These studies show that periodic orbits form if the electron drift momentum vanishes, and assess the role of resonances in stabilization. Vanishing momentum transfer has also been identified as a condition for stabilization using functional analysis methods in a quantum framework [42, 43, 44]. Nonetheless, comparatively little attention has been paid to characterizing the temporal evolution of the KH eigenstates, or coherent superpositions thereof, in phase space. How do these coherent superpositions evolve in time, for a time-averaged or full potential in the KH frame, and what is the physics behind it? To what extent can this evolution be compared with that observed for molecules?
@@
In the present article, we investigate the phase-space dynamics in the KH frame and show that, instead of spatial confinement, stabilization arises from confinement in momentum space. A wave packet constructed as a coherent superposition of KH eigenstates will exhibit a cyclic motion between the two wells of the KH time-averaged dichotomous potential. A phase space analysis employing Wigner quasiprobability distributions in the KH frame shows that, while the position spread of the quasiprobability flow is large, a small spread around vanishing momentum is necessary for stabilization to occur. Stabilization breaks down once this momentum starts to increase, and the full time-dependent dynamics lead to a delay in this cyclic motion. A thorough understanding of these dynamics is important for creating stable light-induced coherent superpositions of states.
@@
This article is organized as follows. In Sec. II, we provide the necessary theoretical background to understand our results and an outline of how our model is implemented. Subsequently, in Sec. III, we perform a series of tests in the key observables to ensure stabilization occurs. In Sec. IV we proceed to a phase-space analysis, first characterizing the dynamics of KH eigenstates and then assessing the phase-space behavior associated with coherent superpositions of KH eigenstates. Finally, in Sec. V we state our conclusions.
@@ II.3 Model and numerical implementation
In Fig. 1, panels a) and b), we plot the probability density of the ground and first excited KH eigenstates (blue solid line), respectively. The ground state wave function is even and has a maximum located at , while the first excited state wave function is odd and has extrema located at the wells of the time-averaged KH potential. The associated eigenenergies are and , and are above the maximum energy of the central KH potential barrier, which is . The KH potential minima are located at approximately a.u.
@@ III Proof of concept
In this section, we analyze different observables to determine the validity of the KH regime and whether the KH eigenstates represent a good basis to model the time-dependent dynamics. As a first case study, we consider the propagation of the ground state of the short-range potential [Eq. (20)] under the action of an electric field of the form (21) by solving the full TDSE [Eq. (1)]. Previous studies [20] have shown that in the high-frequency limit, when , the KH approximation should be valid for any intensity of the laser field. For the parameters used in this work, we have . Additionally, our parameters satisfiy the Gavrila-Kaminsky condition [19, 46], , where is the ground state of the KH time-averaged potential. The present proof of concept builds upon the standard tests performed in order to determine whether the KH regime has been reached.
@@
Fig. 2 displays the width of the wave packet and different bound-state populations, which are the most common observables investigated in this context. The top panels show the results after solving the TDSE [Eq. (1)] using the ground state of the bound potential as initial condition. In Fig 2a), we plot the width of the trapped part of the wave packet as a function of time. We use a spatial filter and consider that for the wave packet is trapped, and compute the width using the standard deviation of the distribution of points within this spatial region. Evidence of reaching the KH regime is given by constant width. The KH regime is only valid during the constant part of the electric field, which starts after a two-cycle turn-on and extends up to . The width remains constant, reaching the value of as discussed in [50]. Evidence of the KH regime is given by the time-dependent population of the KH eigenstates [20]. In Fig 2b) we show how the population of the initial ground state of the short-range potential evaluated from Eq. (11) decays rapidly, while the total population of the KH eigenstates, evaluated from Eq (12), summing over the two eigenstates, remains approximately stationary for the entire time the field amplitude is constant, after a sharp decay occurring during the field turn on. In Figs. 2c) and d) we plot the same observables as in panels a) and b) respectively, but considering the ground state of the KH time-averaged potential as an initial condition to solve the TDSE [Eq. (1)]. We can see overall the same behavior. However, the width of the time-dependent wave packet [Fig 2c)] is narrower and the population of the KH eigenstates [blue line in Fig 2d)] is lower compared to its counterpart in Fig 2b).
@@
In Fig. 3, we investigate the electron’s classical displacement and the expectation value of the position operator, compared with the time-dependent probability density in the laboratory and KH frames. Panel a) shows two periods of the electron classical displacement for a time where the field strength is constant. This displacement exhibits the expected oscillatory behavior and follows the field closely. In Fig 3b), we show snapshots of the probability density in the laboratory frame using the first eigenstate as the initial condition in the full TDSE propagation [Eq. (1)], at the times represented by the color points in Fig 3a). We can see how the probability density presents the dichotomous structure typical of the KH regime [50] and also follows the field in a similar way as the displacement . For instance, at times corresponding to a field zero crossing the two peaks are approximately symmetric about the origin, while, for the times matching the field extrema the probability density is strongly skewed to the left or right. Additionally, in Figs. 3c) and d) we show the effect of transforming from the lab to the KH frame applying Eq. (10) to the probability distribution and the average value of the position of the wave packet. The gauge transformation removes the shifts observed in panel c), so that the probability density in the KH frame no longer follows the field, but still exhibits the same shape and dichotomous structure [Fig. 2d)]. Similarly, comparing in the KH and lab frames [Fig. 2d); orange and blue curves, respectively], one sees that transforming to the KH frame removes the field oscillations to a great extent, but keeps the behavior that develops in a longer timespan.
@@ IV.1 Characterization of the KH eigenstates in phase space
Next, we characterize the KH time-averaged potential in phase space. Although a symmetric coherent superposition of KH eigenstates has been shown in [39], one must assess the Wigner quasiprobability distribution associated with each eigenstate and how coherent superpositions thereof evolve. We also investigate how the Wigner functions behave regarding classical constraints such as equienergy curves given by Eq. (19). In previous work, we have found them to provide insight into the dynamics of different quantum pathways [25, 29, 53, 49].
In Figs. 4a) and b), we show the Wigner distribution (17) for the ground and excited eigenstates of the KH time-averaged potential, respectively, together with the separatrix at a.u. and the equienergy curve for a.u. (white and green curves, respectively). The corresponding probability densities are plotted in Fig. 2. Overall, the Wigner functions associated with the stationary states are roughly contained by the outermost equienergy curve in the figure, although there is a position and momentum spread as required by the uncertainty relation. For the ground KH eigenstate, the Wigner quasiprobability distribution is peaked at the origin, and its width extends beyond the wells determined by the separatrix. In contrast, for the excited state, the Wigner function appears to peak at the turning points defined by the separatrix and to extend beyond, roughly following the shape of the external, green equienergy curve. The Wigner function is also negative at the origin, which is evidence of nonclassical behavior. Since these are the eigenstates of the KH time-averaged potential, we expect the probability density to remain stationary.
@@
and localize the initial wave function in the left well of the KH time-averaged potential. The Wigner quasiprobability distribution corresponding to the coherent superposition in Eq. (24) is shown in Fig. 4c) for . This coherent superposition gives a quasiprobability density with a positive peak on the left well, and a weaker, negative peak on the right well. The momentum ranges remain confined by the external equienergy curve.
@@
In Fig. 5a) we show the autocorrelation function evaluated after solving the TDSE Eq. (8) for the time-averaged KH potential using the two eigenstates separately as initial conditions (superimposed orange and blue lines in the figure) and their equally weighted coherent superposition (green line in the figure). The autocorrelation functions change periodically with frequency indicating how the wave packet oscillates periodically between the two wells of the KH time-averaged potential.
Subsequently, we use the coherent superposition of the two eigenstates, given by Eq. (24), and study its time-dependent dynamics solving the TDSE [Eq. (1)] including the full potential and field. We solve the equation in the lab frame and by applying the transformation (10) we analyze the observables and obtain the Wigner quasiprobability distributions in the KH frame. Fig. 5b) shows the autocorrelation function for this study in the lab frame (blue line) and the KH reference frame (orange line). We can see how by transforming to the KH frame the main oscillations obtained in the KH time-averaged potential (green line) are present. However, the frequency of the oscillations is different. This could be due to the contribution of the time-dependent terms of the expansion or transient effects introduced during the turn-on ramp of the laser field. Nonetheless, the main features of the KH regime are present, with the wave packet moving periodically from one well to the other. A periodic behavior is also observed for a field-free diatomic molecule, for much smaller internuclear separation than the distance between the wells of the KH potential [29, 53]. However, the physical mechanisms underpinning this behavior differ, as we discuss next.

### v2

# Quantum beating and cyclic structures in the phase-space dynamics of the Kramers-Henneberger atom

@@ Abstract
We investigate the phase-space dynamics of the Kramers Henneberger (KH) atom solving the time-dependent Schrödinger equation for reduced-dimensionality models and using Wigner quasiprobability distributions. We find that, for the time-averaged KH potential, coherent superpositions of eigenstates perform a cyclic motion confined in momentum space, whose frequency is proportional to the energy difference between the two KH eigenstates. This cyclic motion is also present if the full time dependent dynamics are taken into consideration. However, there are time delays regarding the time-averaged potential, and some tail-shaped spilling of the quasiprobability flow towards higher momentum regions. These tails are signatures of ionization, indicating that, for the potential studied in this work, a small momentum spread is associated with stabilization. A comparison of the quasiprobability flow with classical phase-space constraints shows that, for the KH atom, the momentum must be bounded from above. This is a major difference from a molecule, for which the quasiprobability flow is confined in position space for small internuclear separation. Furthermore, we assess the stability of different propagation strategies and find that the most stable scenario for the full dynamics is obtained if the system is initially prepared in the KH ground state.
@@ I Introduction
Thus, a legitimate question is how far the analogy between the KH atom and a molecule can be pushed, and how stable the light-induced potentials are. To perform these studies, it is convenient to write the full Hamiltonian in an oscillatory frame of reference, which quivers with the electron in the presence of the laser field: the KH frame. In the KH frame, the time dependence of the field is embedded in the binding potential. Thereby, an important indicator is how far the system’s dynamics are from that obtained for the time-averaged KH potential. This time-averaged potential is the zeroth term of the Fourier expansion of the time-dependent potential and has a dichotomous structure and eigenstates of well-defined shapes and parities. A stable population in the KH eigenstates is evidence that the stabilization regime has been reached [19, 20], as well as the electronic wave packet developing a double-peaked structure. Early studies using Floquet theory also indicate that the Floquet quasienergy states exhibit a structure resembling those of the time-averaged KH potential [21]. Different coherent superpositions of KH eigenstates can be obtained, for instance, by employing fields of different lengths [22].
@@
Important tools for understanding a system’s dynamics are provided by the phase space (for a review see [23]). Besides providing information about the spatial localization of a time-dependent wave packet, phase-space tools also shed light on how its momentum distribution evolves in time, and whether its temporal evolution exhibits nonclassical features. For instance, subtleties in the dynamics of strong-field ionization in atoms [24, 25, 26] and molecules [27, 28, 29, 30], non-dipole time delays in tunneling ionization [31] or high-order harmonic generation in inhomogeneous fields [32], and rescattering [33, 34, 35], and how quantum effects facilitate enhanced ionization [30, 23] have been understood with the help of the phase space. Furthermore, Wigner quasiprobability distributions have been employed to characterize non-classical light [36], and in laser-induced core dynamics [37, 38] and attosecond interferometry [39].
@@
Within strong-field atomic stabilization, quantum phase-space tools have been employed in early work, following the line of inquiry of whether stabilization was classical or quantum [21, 40, 41, 22], and to understand the signatures of wave packets driven by different field types [22]. Tails in the Wigner quasi probability distributions were associated with ionization and a dichotomous structure was linked to the KH potential [41, 22]. Another area of investigation was the nonclassicality of time-dependent terms beyond the zeroth order, associating multiphoton transitions with the corrections to the KH potential [40], time-averaged potentials and scarring in the phase-space distributions [21]. In addition to those, there are also studies using the classical phase space, employing periodic orbit theory and Poincaré sections [42, 43]. These studies show that periodic orbits form if the electron drift momentum vanishes, and assess the role of resonances in stabilization. Vanishing momentum transfer has also been identified as a condition for stabilization using functional analysis methods in a quantum framework [44, 45, 46]. Nonetheless, comparatively little attention has been paid to characterizing the temporal evolution of the KH eigenstates, or coherent superpositions thereof, in phase space. How do these coherent superpositions evolve in time, for a time-averaged or full potential in the KH frame, and what is the physics behind it? To what extent can this evolution be compared with that observed for molecules? Thereby, a key issue is that, if there is population trapping in a molecule, the quasiprobability flow undergoes a cyclic motion, in which direct transfer from one center to the other is observed [27, 28, 29, 30]. This flow is non-classical and associated with quantum interference, which facilitates this transfer. Another pre-requisite for this cyclic motion is spatial confinement, which in phase space is characterized by nested separatrices [30]. The frequency of this cyclic motion is determined by the molecular parameters [47], and can be controlled by an appropriate choice of the driving field parameters [48].
@@
In the present article, we investigate the phase-space dynamics in the KH frame and show that the same cyclic behavior is present for the KH atom if the stabilization regime has been reached. However, a key difference from a molecule is the lack of spatial confinement. A wave packet constructed as a coherent superposition of KH eigenstates will exhibit a cyclic motion between the two wells of the KH time-averaged dichotomous potential. A phase-space analysis employing Wigner quasiprobability distributions in the KH frame for a short-range potential shows that, while the position spread of the quasiprobability flow is large, a small spread around vanishing momentum is observed in the regime for which stabilization occurs. Stabilization breaks down once this momentum starts to increase, and the full time-dependent dynamics lead to a delay in this cyclic motion. A thorough understanding of these dynamics is important for creating stable light-induced coherent superpositions of states.
@@
This article is organized as follows. In Sec. II, we provide the necessary theoretical background to understand our results and an outline of how our model is implemented. Subsequently, in Sec. III, we perform a series of tests in the key observables to ensure stabilization occurs. In Sec. 4 we proceed to a phase-space analysis, first characterizing the dynamics of KH eigenstates and then assessing the phase-space behavior associated with coherent superpositions of KH eigenstates. Finally, in Sec. V we state our conclusions.
@@ II.3 Model and numerical implementation
In Fig. 1, panels (a) and (b), we plot the probability density of the ground and first excited KH eigenstates (blue solid line), respectively. The ground state wave function is even and has a maximum located at , while the first excited state wave function is odd and has extrema located at the wells of the time-averaged KH potential. The associated eigenenergies are and , and are above the maximum energy of the central KH potential barrier, which is . The KH potential minima are located at approximately a.u. The parities of these eigenstates and the shapes associated with the corresponding probability densities resemble those observed in a diatomic molecule. The ground KH eigenstate has even parity and the associated probability density is mainly localized at the midpoint between both wells, while the excited KH eigenstate has odd parity, with the corresponding probability density peaked at the wells.
@@ III Proof of concept
In this section, we analyze different observables to determine the validity of the KH regime and whether the KH eigenstates represent a good basis to model the time-dependent dynamics. As a first case study, we consider the propagation of the ground state of the short-range potential [Eq. (20)] under the action of an electric field of the form (21) by solving the full TDSE [Eq. (1)]. Previous studies [20] have shown that in the high-frequency limit, when , the KH approximation should be valid for any intensity of the laser field. For the parameters used in this work, where and , thus satisfying the condition . Additionally, our parameters satisfiy the Gavrila-Kaminsky condition [19, 50], , where is the ground state of the KH time-averaged potential. The present proof of concept builds upon the standard tests performed in order to determine whether the KH regime has been reached.
@@
Fig. 2 displays the width of the wave packet and different bound-state populations, which are the most common observables investigated in this context. The top panels show the results after solving the TDSE [Eq. (1)] using the ground state of the bound potential as initial condition. In Fig 2(a), we plot the width of the trapped part of the wave packet as a function of time. We use a spatial filter and consider that for the wave packet is trapped, and compute the width defined as twice the standard deviation of the distribution of points within this spatial region. Evidence of reaching the KH regime is given by constant width. The KH regime is only valid during the constant part of the electric field, which starts after a two-cycle turn-on and extends up to . The width remains constant, reaching the value of as discussed in [53]. Further evidence of the KH regime is given by the time-dependent population of the KH eigenstates [20]. In Fig 2(b) we show how the population of the initial ground state of the short-range potential evaluated from Eq. (11) decays rapidly, while the total population of the KH eigenstates, evaluated from Eq (12), summing over the two eigenstates, remains approximately stationary for the entire time the field amplitude is constant, after a sharp decay occurring during the field turn on. In Figs. 2(c) and (d) we plot the same observables as in panels (a) and (b), respectively, but considering the ground state of the KH time-averaged potential as an initial condition to solve the TDSE [Eq. (1)]. We can see the same overall behavior. However, the width of the time-dependent wave packet [Fig 2(c)] is narrower and the population of the KH eigenstates [blue line in Fig 2(d)] is lower compared to its counterpart in Fig 2(b).
In Fig. 3, we investigate the electron’s classical displacement and the expectation value of the position operator, compared with the time-dependent probability density in the laboratory and KH frames. Because we are interested in the part of the wave packet trapped by the KH potential, the expectation value of the position operator is calculated for the same region in space, that is, . It is worth noticing that within this region the probability density drops before the onset of stabilization due to ionization. Hence, to define the expectation value we need to enforce normalization of the trapped part of the wave packet to make up for irreversible ionization. The expectation value of the position operator is then defined as
@@
Fig. 3(a) shows two periods of the electron classical displacement for a time where the field strength is constant. This displacement exhibits the expected oscillatory behavior and follows the field closely. In Fig 3(b), we show snapshots of the probability density in the laboratory frame using the first eigenstate as the initial condition in the full TDSE propagation [Eq. (1)], at the times represented by the color points in Fig 3(a). We can see how the probability density exhibits the dichotomous structure typical of the KH regime [53] and also follows the field in a way similar to the displacement . For instance, at times corresponding to a field zero crossing [see the blue, green and purple dots in Fig. 3(a)], the two peaks in the corresponding probability densities [blue, green and purple curves in Fig. 3(b)] are approximately symmetric about the origin, while, for the times matching the field extrema [yellow and red dots in Fig. 3(a)] the probability density is strongly skewed to the left or right [see yellow and red curves in Fig. 3(b)]. In Fig. 3(c) we plot the expectation value of the position operator as defined in Eq. (24). The expectation value of the trapped part of the wave packet oscillates approximately between following the field oscillations in the lab frame (blue line). Additionally, in Figs. 3(c) and (d) we show the effect of transforming from the lab to the KH frame applying Eq. (10) to the probability distribution and the average value of the position of the wave packet. The gauge transformation removes the shifts observed in panel (b), so that the probability density in the KH frame no longer follows the field, but still exhibits the same shape and dichotomous structure [Fig. 2(d)]. Similarly, comparing in the KH and lab frames [Fig. 2(c); orange and blue curves, respectively], one sees that transforming to the KH frame removes the field oscillations to a great extent, but keeps the behavior that develops in a longer timespan.
@@ IV Phase-space analysis and stability
The Wigner quasiprobability distributions [Eq. (17)] mirror the behavior observed in Figs. 2 and 3, showing evidence of stabilization. In Fig. 4, we display the Wigner quasiprobability distributions calculated assuming that the initial state is the ground state of the atomic potential, as shown in Eq. (20). These quasiprobability densities are in the KH reference frame. The times have been selected by analyzing Fig. 2(b) to find relevant moments of the dynamics, aiming to demonstrate the onset of the KH atom in phase space. The initial Wigner distribution, shown in Fig. 4(a) for , is centered at the origin, as expected for the atomic ground state. Subsequently, Fig. 4(b) shows the Wigner function at a.u., which corresponds to a transient region where stabilization has not set in. The original shape of the quasiprobability density is distorted and strong fringes are present around it. Figs. 4(c) and (d) show the quasiprobability flow at a.u. and a.u., respectively, which is after the onset of stabilization. This onset is indicated by the emergence of the dichotomous structure in the probability density which led to the observed constant width of the trapped wavepacket shown in Fig. 2. The Wigner function elucidates the constant width of the wavepacket in phase space and appears to exhibit a cyclic behavior around the separatrix. The flow associated with this cyclic behavior stays within the green equienergy curve. There are also strong fringes/tails escaping the boundaries of the equienergy curve. The quasiprobabilities in panels (c) and (d) are very similar except for a reflection. In order to understand this behavior better, we must inspect the KH atom in phase space.
@@ IV.1 Characterization of the KH eigenstates in phase space
Next, we characterize the KH time-averaged potential in phase space. Although a symmetric coherent superposition of KH eigenstates has been shown in Ref. [41], one must assess the Wigner quasiprobability distribution associated with each eigenstate and how coherent superpositions thereof evolve. We also investigate how the Wigner functions behave regarding classical constraints such as equienergy curves given by Eq. (19). In previous work, we have found them to provide insight into the dynamics of different quantum pathways [25, 30, 47, 48].
In Figs. 5(a) and (b), we show the Wigner distribution (17) for the ground and excited eigenstates of the KH time-averaged potential, respectively, together with the separatrix at a.u. and the equienergy curve for a.u. (white and green curves, respectively). The corresponding probability densities are plotted in Fig. 3. Overall, the Wigner functions associated with the stationary states are roughly contained by the outermost equienergy curve in the figure, although there is a position and momentum spread as required by the uncertainty relation. For the ground KH eigenstate [Fig. 5(a)], the Wigner quasiprobability distribution is peaked at the origin, and its width extends beyond the wells determined by the separatrix. In contrast, for the excited KH eigenstate [Fig. 5(b)], the Wigner function appears to peak at the turning points defined by the separatrix and to extend beyond, roughly following the shape of the external, green equienergy curve. The Wigner function is also negative at the origin, which is evidence of nonclassical behavior. Since these are the eigenstates of the KH time-averaged potential, we expect the probability density to remain stationary.
@@
and localize the initial wave function in the left well of the KH time-averaged potential. The Wigner quasiprobability distribution corresponding to the coherent superposition in Eq. (25) is shown in Fig. 5(c) for . This coherent superposition gives a quasiprobability density with a positive peak on the left well, and a weaker, negative peak on the right well. The momentum ranges remain confined by the external equienergy curve. A similar shape is observed for a diatomic molecule, if a localized coherent superposition of the ground and excited eigenstates is constructed [30].
@@
In Fig. 6a) we show the absolute value of the autocorrelation function defined as
evaluated after solving the TDSE Eq. (8) for the time-averaged KH potential using the two eigenstates separately as initial conditions (superimposed dashed gray and blue lines in the figure) and their equally weighted coherent superposition as defined in Eq. (25) (green line in the figure). The autocorrelation functions change periodically with frequency indicating how the wave packet oscillates periodically between the two wells of the KH time-averaged potential.
Subsequently, we study the time-dependent dynamics of the initial state given by the coherent superposition in Eq. (25) considering the full potential and electric field. We solve the TDSE Eq. (1) in the lab frame and by applying the transformation (10) we analyze the observables and obtain the Wigner quasiprobability distributions in the KH frame. Fig. 6(b) shows the absolute values of the autocorrelation function for this study in the lab frame (blue line) and the KH reference frame (orange line). As a reference we have included the calculations done for the KH time-averaged potential shown with the green lines in both panels in Fig. 6. By comparing the results in the KH frame (orange line) with those computed using the KH time-averaged potential (green line), we can see how by transforming to the KH frame the main oscillations obtained in the KH time-averaged potential, which are associated with the cyclic behavior of the Wigner functions, are present. However, the frequency of the oscillations is different. This could be due to the contribution of the time-dependent terms of the expansion or transient effects introduced during the turn-on ramp of the laser field. Nonetheless, the main features of the KH regime are present, with the wave packet moving periodically from one well to the other. A periodic behavior is also observed for a field-free diatomic molecule, for much smaller internuclear separation than the distance between the wells of the KH potential [30, 47]. However, the physical mechanisms underpinning this behavior differ, as we discuss next.